### v3

@@ 3.2.1 StreamingLLM Ring Buffer
In this section, we motivate and describe the implementation of the StreamingLLM rolling KV cache window, which is implement with a ring buffer. For a naïve implementation of StreamingLLM using a static KV cache size , generating beyond tokens one at a time involves inefficiently slicing the sink tokens, recent token window, and top- tokens from the KV cache and concatenating them all together with the new keys and values (see L19-26 of Listing 2 in the Appendix). When the rolling window is implemented as a ring buffer, we only need to scatter new keys and values into the appropriate index during decoding (see Figure 3(d)). This requires us to prepare the ring buffer structure during the prefill phase, which involves two gather calls for the sequence indices before and after the prefill length index. See Listing 4 in the Appendix for a PyTorch-style implementation of the ring buffer construction, and Figure 4 for a visualization of where the two gather calls are placed along the sequence dimension.
@@ Appendix A SnapStream Static Graph Pseudo-code
In this section, we show and describe PyTorch-style pseudocode for the operation of a SnapStream KV cache during prefill and decoding. We compare the implementations of a naïve implementation of SnapStream with a rolling recent window buffer, and our instantiation with a ring buffer.
@@ A.1 Rolling Recent Window
Listing 1 shows an implementation of the KV cache compression process for a naïve implementation of SnapStream with a rolling window buffer, as in StreamingLLM. It uses the prefill_snapstream_topk method described in Listing 6 to extract the top- evicted tokens between the sink and recent tokens. Construction of the compressed KV cache is then a straightforward concatenation in Lines 22-23.
Listing 2 shows an implementation of the decoding phase KV cache update using the naïve implementation of SnapStream with a straightforward recent window buffer. In the case where the cache_position (or sequence length in previous sections) is less than max_kv_length, the update is a straightforward scatter (Lines 15-16). In the other case (Lines 19-26), we construct a new KV cache by concatenating slices of the sink, Top-, and recent KVs. To accommodate different cache_position values at different batch indices in a static graph, we perform both KV cache updates and select between the two depending on the batch-indexed cache position on Line 28. In practice, the implementation of Branch 2 in this code leads to materialization of many intermediate buffers that take up extra on-chip memory. For dataflow architectures in particular, this can lead to inefficient allocation of pattern memory units during placing and routing that prolong program load and increase latency.
@@ A.2 Ring Buffer
Then, we use prefill_streamingllm_slicing to construct the ring buffer for the full SnapStream KV Cache in Listing 4. This greatly simplifies the decoding stage to a single scatter for each of the key and value tensors in Listing 5.

### v4

@@ 3.2.1 StreamingLLM Ring Buffer
In this section, we motivate and describe the implementation of the StreamingLLM rolling KV cache window, which is implemented as a ring buffer. For a naïve implementation of StreamingLLM using a static KV cache size , generating beyond tokens one at a time involves inefficient slicing of the sink tokens, recent token window, and top- tokens from the KV cache and concatenating them all together with the new keys and values (see Lines 24-32 of Listing 2 in the Appendix). When the rolling window is implemented as a ring buffer, we only need to scatter new keys and values into the appropriate index during decoding (see Figure 3(d)). This requires us to prepare the ring buffer structure during the prefill phase, which involves two gather calls for the sequence indices before and after the prefill length index. See Listing 3 in the Appendix for a PyTorch-style implementation of the ring buffer construction, and Figure 4 for a visualization of where the two gather calls are placed along the sequence dimension.
@@ Appendix A SnapStream Static Graph Pseudo-code
In this section, we show and describe PyTorch-style pseudocode for the operation of a SnapStream KV cache during prefill and decoding. We compare a naïve implementation of SnapStream with a rolling recent window buffer, and our instantiation with a ring buffer.
@@ A.1 Rolling Recent Window
Listing 1 shows an implementation of the KV cache compression process for a naïve implementation of SnapStream with a rolling window buffer, as in StreamingLLM. It uses the prefill_snapstream_topk method described in Listing 6 to extract the top- evicted tokens between the sink and recent tokens. Construction of the compressed KV cache is then a straightforward concatenation in Lines 25-26.
Listing 2 shows an implementation of the decoding phase KV cache update using the naïve implementation of SnapStream with a rolling recent window buffer. In the case where cache_position (sequence length in previous sections) is less than max_kv_length, the update is a straightforward scatter (Lines 19-22). In the other case (Lines 24-32), we construct a new KV cache by concatenating slices of the sink, Top-, and recent KVs. To accommodate different cache_position values at different batch indices in a static graph, we perform both KV cache updates and select between the two depending on the batch-indexed cache position on Line 34. In practice, the implementation of case 2 in this code snippet leads to materialization of many intermediate buffers that take up extra on-chip memory. For dataflow architectures in particular, this can lead to inefficient allocation of pattern memory units during placing and routing that prolong program load and increase latency.
@@ A.2 Ring Buffer
Then, we use prefill_streamingllm_slicing to construct the ring buffer for the full SnapStream KV Cache in Listing 4. This greatly simplifies the decoding stage to a single scatter for each of the key and value tensors in Listing 5. Note that in the pseudocode, Range 1 and Range 2 are swapped when compared to Figure 4.